OpenStudy (anonymous):

MEDAL FOR BEST ANSWER What is the percent of decrease from 110 to 77? A. 30% B. 33% C. 43% D. 70%

OpenStudy (anonymous):

answer is A.30%

OpenStudy (anonymous):

SO, let us begin with 77 is what percent of 110? 110 (x) = 77 (110 times what decimal form percentage yields 77?) 110 (x) / 110 = 77 / 110 (Divide both sides by 110; simplify.) x = 0.70 (Or in percentage form 70%; therefore, 77 is 70% of 110. Now, we can subtract this from 100% to determine the exact decrease that took place to reach 77 from 110.) 100% - 70% = 30% (Therefore, our percent decrease is exactly a 30% decrease from 110 to 77.)

OpenStudy (anonymous):

THANKS

OpenStudy (anonymous):

the answer is 70%

OpenStudy (anonymous):

the reason I say that is if you multiply 110 by 0.7 or 70% toy get 77